OpenStudy (anonymous):

Mark and Nancy are making punch for a party and are using two types of juice: pineapple and papaya. The total number of quarts must be more than 9 but less than 15. They should use at least 4 quarts of each type. We want to find out how many quarts of each juice they should buy. List any three ordered pair solutions from your graph in #1.

OpenStudy (anonymous):

Okay. We get 4 equations, right? x>=4 y>=4 x+y>=9 and x+y <= 15 To find vertices, you just need to pic two of those equations and figure out if (and where) they cross.

OpenStudy (anonymous):

yeah...can u just give me the ordered pairs?

OpenStudy (anonymous):

Nope. I don't roll like that.

OpenStudy (anonymous):

aww

OpenStudy (anonymous):

please! this 1 time?!

OpenStudy (anonymous):

Well x+y >= 9 and x+y <=15 don't cross because the two lines have the same slope. -1. So you need to get your ordered pairs by matching up the other equations. Use x=4 and y = -x+9 to get one ordered pair. Use y = 4 and y = -x+15 to get another. Use x = 4 and y = 4 to get another.

OpenStudy (anonymous):

yay! thanks!